OpenStudy (anonymous):

A right triangle has a hypotenuse of 153 centimeters and a leg of 72 centimeters. What is the length of the other leg?

OpenStudy (anonymous):

\[a ^{2} + b ^{2} = c ^{2}\] where 'a' is one of the 'legs' and 'c' is the hypotenuse.\[a = 72, c = 153\] the problem then becomes \[5184 + b ^{2} = 23409\] \[b ^{2} = 23409 - 5184 = 18225\]\[b = \sqrt{18225} = 135\]